OpenStudy (anonymous):

can u guys helpp?? Solve for x: 3(x + 1) = -2(x - 1) + 6.

OpenStudy (chelseatheruler):

3x+3=-2x+2+6 5x+3=8 5x=5 x=1

OpenStudy (chelseatheruler):

you're welcome.

OpenStudy (anonymous):

thank you so much

OpenStudy (chelseatheruler):

Np. Medal?

OpenStudy (anonymous):

distribute 3 unto x+1 so it would be 3x+3 then do the same thing on the other side, -2 unto x-1 so it would be -2x+2+6 then gather like terms so it would be 3x+3=-2x+8, subtract the 3 unto the 8 and add the -2x unto the 3x. so it would look like this 5x=5 divide by the 5 and get x alone to equal 1

OpenStudy (anonymous):

so the answer would be 1 right?

OpenStudy (anonymous):

yes x=1